OpenStudy (anonymous):

what is a cubic polynomial functions in standard form with zeros 3, -6, and 0

OpenStudy (jdoe0001):

\(\bf x=3\implies {\color{blue}{ x-3}}=0\quad x=-6\implies {\color{blue}{ x+6}}=0\quad x=0\implies {\color{blue}{ x}}=0\\ \quad \\ (x-3)(x+6)(x)=0\implies (x-3)(x+6)(x)=\textit{original cubic equation}\\ \quad \\ \implies (x-3)(x+6)(x)=y\) solve for "y" to make it in standard form